### v2

# A First Look at Hydrogen Generation in an Ultramafic Rock with Micro-CT and SEM-BEX

Abstract
Natural hydrogen (H2) generated by water–rock interaction in ultramafic rocks is increasingly recognised as a potentially important primary energy resource, but the pore-scale processes that control the initiation and early transport of a free gas phase remain poorly constrained. Here we present an in situ X-ray micro-tomography experiment in which an ultramafic granular pack of ultramafic from Papua Province, Indonesia, saturated with KI-doped brine, is heated to 100C with a pore pressure of 4 bar under 10 bar confining pressure inside a micro-CT scanner. Time-resolved 4D imaging captures the transition from a fully liquid-saturated pore space to the appearance and growth of a distinct gas phase after an $\sim$8 h induction period. Bubbles first nucleate near the top of the sample before becoming distributed throughout the imaged volume as a connected ganglia. The nucleating gas phase is most plausibly dominated by molecular hydrogen (H2) generated by low-temperature fluid–rock reaction, as indicated by independent hydrogen-presence detectors, although we cannot yet fully exclude minor contributions from other gases. SEM-BEX imaging reveals textural alteration and local changes in elemental signals between reacted and unreacted material. Taken together, these observations provide spatially and temporally resolved evidence for gas generation during low-temperature alteration of ultramafic grains and demonstrate that pore-scale imaging can directly link water–rock reaction kinetics, gas generation and multiphase flow behaviour in natural hydrogen systems.
[…]
Institute of GeoEnergy Engineering, Heriot-Watt University, Edinburgh, UK The Lyell Centre, Heriot-Watt University, Edinburgh, UK
[…]
*Hannah P. Menkeh.menke@hw.ac.uk
{keypoints}
We use 4D X ray micro-CT to image gas formation in a reacting ultramafic rock
A gas phase appears during low temperature reaction of brine saturated ultramafic grains
Electron microscopy shows mineral alteration and element redistribution consistent with fluid rock reaction
Observations provide qualitative pore scale evidence relevant to natural hydrogen generation in the subsurface
Plain Language Summary
Hydrogen gas can be produced naturally when certain ultramafic, iron-rich rocks formed in Earth’s mantle react with water. There is growing interest in whether this “natural hydrogen” could be used as a low-carbon energy source, but we still know very little about where the gas forms inside the rock and how it starts to move.
In this study, we packed grains of an ultramafic rock into a small cylinder, saturated them with salty water, and heated the system while it sat inside a 3D X-ray scanner (micro-CT). This let us watch the pore space inside the sample over time. At first, all of the pores contained only liquid. As the experiment continued, dark patches appeared and grew in the images, showing that a separate gas phase was forming. Two independent measurements of the gas produced during the experiment are consistent with hydrogen, although this single experiment cannot fully describe all of the reactions taking place.
After imaging, we examined the reacted grains under an electron microscope and saw evidence that the minerals had altered and some elements had moved. Together, these observations demonstrate that we can directly visualise gas generation inside reactive rocks at relatively low temperatures. This is an early, exploratory step towards understanding how natural hydrogen might form and behave in the subsurface.
1 Introduction
Natural geologic hydrogen (H2) is emerging as a potentially vast, carbon-free primary energy resource that could complement, or in some settings substitute, engineered “green” H2 production. Global compilations of gas analyses from seeps, boreholes and mines show that H2 is far more widespread than previously appreciated and occurs in many tectonic settings, often as a major component of the gas phase (Zgonnik2020). Recent system-scale syntheses formalise the concept of a “H2 system”, in which H2 is generated, migrates and accumulates in traps in an analogous way to hydrocarbons (Jackson2024; Prinzhofer2015), and multi-scale reviews highlight the key migration pathways and timescales from the planetary to basin scale (Lodhia2024_H2Migration). This has led to growing industrial interest in “native” or “natural” hydrogen, with first-order assessments suggesting that alteration-driven systems (e.g. areas of serpentinisation) could sustain commercially significant fluxes over geological timescales (Gaucher2020), and to proposals for engineered “rock-based” hydrogen concepts (white and “orange” hydrogen) that explicitly exploit water–rock reactions in mafic and ultramafic rocks as part of the low-carbon hydrogen portfolio (Osselin2022).
Among the diverse proposed sources of natural H2, low- to high-temperature serpentinization of ultramafic rocks is generally considered the most efficient and volumetrically important (Oze2007; McCollom2009; Jackson2024). In this process, Fe2+-bearing primary minerals such as olivine and pyroxene react with water to form Fe3+-bearing phases (e.g., magnetite, Fe(III)-serpentine), generating H2 as a by-product (Sleep2004; Holm2015). Thermodynamic and kinetic models predict that H2 yields depend sensitively on temperature, water-to-rock ratio, mineralogy, water chemistry, and redox state (McCollom2009; McCollom2016; Ely2023), while field observations from ophiolites and oceanic and continental settings (e.g., Samail ophiolite, Oman; Lost City hydrothermal field) demonstrate that serpentinization can sustain long-lived fluxes of H2-rich, hyperalkaline fluids (Schrenk2013). More recently, work in the Samail ophiolite Oman has shown that significant H2 can also be generated during late, low-temperature ($200^{\circ}$C) alteration of already serpentinized peridotite, extending the window of H2 production to cooler, near-surface conditions (Ellison2021).
[…]
Core-scale reactive transport experiments bridge part of this gap by preserving rock fabric and permeability while allowing continuous geochemical monitoring. Percolation of reactive fluids through ultramafic rocks and serpentinites has demonstrated strong couplings between serpentinization, carbonation, permeability reduction and H2 production (e.g., Wang2019; Albers2021; Farough2016). Yet in most such studies the rock behaves as a “black box”: H2 production is inferred from outlet fluid chemistry or headspace analyses, and any gas phase that forms inside the pore network remains invisible. Even in well-characterised natural systems such as the Samail ophiolite, constraints on H2 generation and transport rely primarily on fluid chemistry, isotopes and mineralogical proxies rather than direct imaging of the gas phase itself (Miller2016; Ellison2021). As a result, the fundamental pore-scale mechanisms governing the initiation and early transport of H2, bubble nucleation sites, snap-off events, critical gas saturation and connectivity, are still poorly constrained.
In parallel, a large and rapidly growing body of work has used X-ray micro-tomography (micro-CT) to image dynamic multiphase flow and reactive transport in rocks. Time-resolved (4D) micro-CT has been applied to dissolution and precipitation in carbonates and sandstones, enabling direct observations of evolving pore structures, wormholing and reaction-front dynamics under reservoir conditions (NoirielSoulaine2021; NoirielRenard2022; Menke2015). Similar techniques have been used to quantify residual trapping and ganglion dynamics during CO2–brine flow, to measure capillary pressure–saturation relationships, and to calibrate pore-network and direct numerical simulations (Andrew2014; Jiang2015; Bultreys2016). These studies demonstrate that 4D imaging can resolve micron-scale fluid–fluid and fluid–solid interfaces, capturing non-linear displacement patterns, snap-off and reconnection events that strongly influence macroscopic transport properties.
Only recently has this pore-scale imaging toolkit been applied to H2 itself. Micro-CT studies of H2–brine flow in sandstones show that H2 behaves as a non-wetting phase in many reservoir-relevant conditions, with significant capillary trapping, strong sensitivity to rock heterogeneity and subtle but important differences from CO2–brine systems (Thaysen2023; Jangda2023; Rezaei2022). These works provide critical benchmarks for underground H2 storage, but they all consider externally injected H2 displacing brine; the H2 is not generated in situ by water–rock reaction. Complementary core-flooding experiments and relative permeability measurements reveal that H2–water flow in rocks is typically capillary-limited, with low minimum water saturations and relative permeability curves that differ from those of CO2–water systems (Yekta2018; Rezaei2022; Boon2022). Reviews of underground H2 storage emphasise that uncertainties in pore-scale trapping, dissolution and reaction with the rock and microbiota are among the key scientific challenges for safe and efficient large-scale deployment (Heinemann2021; Miocic2023).
Despite these advances, a critical observational gap remains at the intersection of serpentinization geochemistry and pore-scale multiphase flow physics. To our knowledge, there are no prior studies that directly image the nucleation and growth of an H2 gas phase as it is produced within an ultramafic rock under hydrothermal conditions. Existing serpentinization experiments constrain bulk H2 generation rates but lack spatially resolved information on gas formation, while pore-scale H2 imaging studies focus on storage scenarios with externally supplied gas. This missing link limits our ability to assess whether H2 generated during low-temperature alteration can reach saturations and connectivity sufficient for migration and accumulation, or whether it is predominantly trapped and dissolved within low-permeability alteration halos.
In this study, we address this gap using a proof-of-concept in situ experiment that combines time-resolved 4D micro-CT imaging with post-mortem SEM and BEX analyses of an ultramafic granular pack reacting with brine at 100C and 4 bar pore pressure. Our approach allows us to (i) document the transition from a fully liquid-saturated pore network to the appearance and subsequent growth of a distinct gas phase, (ii) qualitatively characterise the spatio-temporal patterns of bubble nucleation and coalescence within the rock framework, and (iii) relate these physical observations to textural and compositional changes observed in SEM-BEX images, including the development of secondary phases and local variations in elemental signals. When considered together with independent measurements of gas generation, these observations are consistent with low-temperature H2 production via water–rock reaction, but they do not on their own uniquely fully resolve the reaction pathways. We therefore present this single, exploratory experiment as an initial step towards linking serpentinization-driven gas generation with pore-scale transport physics in natural H2 systems, and we explicitly highlight the limitations and uncertainties that future, more targeted studies will need to address.
2 Materials and Methods
2.1 Sample Selection and Preparation
The experiment utilized an ultramafic rock sample selected for its potential for low-temperature serpentinization. The sample was collected from the mantle sequence of the Cyclops Ophiolite located in the Cyclops Mountains of Papua Province, Indonesia (along the northern coast of the island of New Guinea). Petrographic inspection shows extensive alteration consistent with serpentinization, with relict olivine grains preserved within a fine-grained alteration matrix.
The rock was crushed and dry sieved to isolate the coarse fraction, retaining only grains between 0.6-1 mm. This grain size was selected to ensure a resolvable pore network for 3D imaging while maintaining reactive surface area.
2.2 Experimental Setup and Protocol
All experiments were conducted at a temperature of 100C. The pore space was saturated with KI-doped brine at a pore pressure of 4 bar, while the core holder was loaded within a Hassler-type sleeve under an external confining pressure of 10 bar. Unless otherwise stated, pressures quoted in this study refer to pore pressure. At the experimental temperature of 100C and a pore pressure of 4 bar, the KI-doped brine is well below its saturation vapour pressure (water boils at 100C only at $\sim$1 bar), although a small amount of water vapour will always be present in any gas phase in equilibrium with liquid water. We therefore do not expect boiling or a persistent steam phase under these conditions; any free gas that appears must arise from exsolution of non-condensable species rather than liquid water boiling.
The granular sample was loaded into a custom-built X-ray transparent core holder designed for high-pressure and high-temperature flow experiments. A strip of H2 indicator was attached inside the confining sleeve of the core holder to detect the presence of H2. The rig was mounted within the X-ray micro-tomography (micro-CT) scanner to facilitate in situ imaging (Fig. 1).
The initial saturation protocol was designed to eliminate trapped air and ensure full liquid saturation prior to heating. First, the sample was flushed with CO2 at ambient conditions. The system was then confined at 10 bar pressure. Subsequently, the sample was flooded with a brine solution composed of 5 wt% potassium iodide (KI) at a pore pressure of 4 bar to keep the brine in the liquid phase. KI was used as an X-ray contrast agent to enhance the attenuation of the fluid phase, thereby maximizing the contrast between the rock grains (grey), the brine (bright white), and any generated gas (black). At this concentration KI is not expected to strongly modify the low-temperature alteration reactions compared to chloride brines, and we did not observe any textural or compositional evidence for abundant KI salt precipitates in the post-reaction samples. The system was flooded for 1000 Pore Volumes (PV) to ensure the complete dissolution of the initial CO2 into the brine, ensuring that any gas phase observed later resulted from reaction rather than trapped shielding gas.
Figure 1: Experimental set-up for in situ micro-CT imaging of H2 generation and gas exsolution. A Hassler-type core holder containing crushed ultramafic rock is mounted on the rotation stage inside a Pb-lined micro-CT enclosure. The sample is confined by a Viton sleeve and aluminium foil within an aluminium outer sleeve, with steel end caps and end fittings connecting to the flow lines. Heating tape and a thermocouple control and monitor the sample temperature, and H2-indicator tape is placed around the core to detect any leakage. Pore fluid is injected from the injection pump, while the receiving pump maintains pressure and collects effluent. Confining pressure is applied via the confining pump. The red dashed box indicates the imaged region within the core.
Following the brine saturation phase, flow was halted (static conditions with no net through-flow), and the temperature of the sample was raised to 100C to initiate the hydrothermal reaction. During this static stage, pore pressure is maintained at 4 bar, so any volumetric expansion associated with gas exsolution was accommodated by small displacements of brine into the pump rather than by a significant increase in pore pressure.
Following the conclusion of the in situ experiment, the system was degassed and the H2 concentration in the pump gas effluent was monitored using a portable electrochemical H2 detector (RAE ToxiRAE Pro PGM-1860). This instrument provides presence/threshold measurements of H2 in the ppm range rather than a full gas composition, but it allows us to confirm whether H2 is present above background levels during controlled depressurisation. The reacted sample was then removed from the core holder and the H2-indicator strip was retrieved from the confining sleeve.
2.3 Micro-CT Image Acquisition
[…]
2.4 Micro-CT Image Analysis
To quantify the evolution of the gas phase, all time-lapse micro-CT volumes were processed using a consistent image-analysis workflow in the image processing software Avizo 2025.1 (Thermo Fisher Scientific). First, each scan was rigidly registered to the initial reference scan by maximising mutual information between grey-scale volumes and resampled onto the reference grid using a Lanczos kernel (Lanczos1956) to preserve edge sharpness. Second, we applied a non-local means filter to reduce high-frequency noise while retaining phase boundaries in the greyscale images (Buades2005). Third, the solid, brine and gas phases were segmented in the Avizo Segmentation Editor using simple global grey-level thresholding, guided by the histogram and visual inspection of representative slices. Conservative thresholds were chosen to separate solid from pore space, and within the pore space voxels with the lowest attenuation were classified as gas. Small isolated gas clusters below a few voxels were removed as likely noise using basic morphological operations. A representative 2D slice illustrating the raw, filtered and segmented images is shown in Figure 2. For each time step we then calculated bulk gas saturation as the ratio of gas-filled voxels to pore-space voxels. Visual inspection of representative slices and 3D renderings was used to check the segmentation quality; we therefore regard the resulting saturation values as approximate but internally consistent across the time series.
Figure 2: Representative 2D slice showing the micro-CT image processing workflow. (a) Raw greyscale image, (b) the same slice after non-local means filtering to reduce high-frequency noise while preserving phase boundaries, and (c) final three-phase segmentation into solid grains (dark blue), brine-filled pore space (red) and gas voxels (light blue).
To check that our conclusions do not depend on a single arbitrary choice of thresholds, we also inspected segmentations generated using slightly more and less conservative gas–brine thresholds that were still consistent with the grey-level histograms and visual appearance of the images. These alternative segmentations yielded very similar temporal trends in bulk $S_{g}$ and connectivity, and in particular preserved the three-stage evolution (no resolvable gas, rapid increase, slower late-time growth).
Gas ganglia statistics were obtained from the segmented gas phase using Avizo’s Label Analysis module. Connected-component labelling was applied to the binary gas mask to identify individual gas clusters (ganglia), and for each cluster we extracted its volume. From these we computed the distribution of ganglia sizes, the number density of clusters, and the volume fraction of the largest cluster $f_{\max}$ at each time step.
Percolation behaviour was assessed by checking whether the largest cluster formed a continuous path spanning the sample between opposing faces (top–bottom and/or side–side) based on voxel-wise connectivity. In other words, a cluster was classified as percolating if it provided a continuous, spanning pathway through the imaged volume; large, internally connected ganglia that did not satisfy this spanning criterion were treated as non-percolating.
2.5 SEM and BEX analysis
Pre- and post-reaction grains for electron microscopy were prepared from the same bulk material as the micro-CT sample. Sub-samples were gently dry-sieved to remove dust, sprinkled onto aluminium stubs, and immobilised using conductive adhesive carbon tabs (Agar Scientific). No specific rinsing step was applied to remove KI brine residues prior to drying, so we cannot completely exclude the possibility that a minor fraction of the observed surface features reflects salt precipitation upon drying. However, the platy surface crystals and roughened patches we focus on occur in discrete areas whose BEX signatures are dominated by Mg, Al, Ca and Fe rather than K or I (Table 2), and their morphology differs from simple euhedral halide crystals. We therefore interpret them primarily as alteration-related features rather than dried KI.
The mounts were imaged using an FEI Quanta FEG 650 Scanning Electron Microscope (SEM) operated in low-vacuum mode. Backscattered electron (BSE) images were acquired at magnifications of hundreds to several thousand times to capture both grain-scale morphology and fine-scale surface textures. For each mount, we collected combined backscatter and energy-dispersive X-ray (BEX) data using an Oxford Instruments Unity detector in combination with an X-MaxN 150 mm Energy Dispersive X-ray (EDX) detector. These generate colour-composite elemental maps by assigning selected element signals (e.g., Mg, Al, Si, Ca, Fe) to different colour channels along with quantitative elemental analysis. These BEX composites were used to visualise spatial variations in surface composition across many grains using the Oxford Instruments AZtec 6.2™ software. In addition, five spot elemental analyses were acquired on each mount, targeting representative areas on the larger grains, and the resulting ranges in composition are reported in Table 2. No matrix-specific corrections were applied beyond the standard instrument routines, so the spot analyses are interpreted qualitatively and in conjunction with the textural information from the BSE and BEX images.
3 Results
3.1 In situ observation of H2 nucleation and growth
The 4D micro-CT scans provided a direct visualization of gas phase evolution within the pore space. During the initial heating and early static reaction period, the pore space remained fully brine-saturated (Figure 3a), with an image-derived porosity of $\phi=0.375$ and a specific solid–pore interfacial area of $A/V_{s}=1.13\times 10^{4}\ \mathrm{m^{2}\,m^{-3}}$ (Table 1a). No gas voxels were detected at this stage ($S_{g}=0$).
Figure 3: Time series of segmented gas phase (false-colour red) within the ultramafic granular pack at (a) 0, (b) 8, (c) 23, and (d) 30 hours. The grains are rendered as translucent grey and the brine is transparent, so that the red overlay highlights only the voxels classified as gas in the micro-CT segmentation.
Gas bubble nucleation was first observed at 8 hours, when a small number of discrete bubbles appeared, localised primarily at the top of the sample column (Figure 3b). At this time the bulk gas saturation remained very low ($S_{g}\approx 0.01$; Table 1b), indicating that only a tiny fraction of the pore space had exsolved. Importantly, the experiment had already been at (approximately) constant temperature and pore pressure for several hours by this point, so this delayed onset is difficult to reconcile with simple thermal expansion of trapped air or immediate exsolution of pre-existing dissolved gas during the initial heating ramp. Instead, it is more naturally interpreted as the time required to generate and accumulate reaction-derived H2 to supersaturation.
By the subsequent scan at 23 hours, bubble nucleation and growth had propagated throughout the imaged volume (Figure 3c). The bulk gas saturation had increased to $S_{g}\approx 0.39$, corresponding to an average rate of change $\Delta S_{g}/\Delta t\approx 2.6\times 10^{-2}\ \mathrm{h^{-1}}$ between 8 and 23 hours (Table 1b; Figure 5). This period therefore marks a rapid onset of widespread supersaturation and gas generation within the pore space. Continued observation at 30 hours showed a further but more modest increase in bulk gas saturation to $S_{g}\approx 0.43$, with a much lower apparent rate of change ($\Delta S_{g}/\Delta t\approx 5.9\times 10^{-3}\ \mathrm{h^{-1}}$) over the final 7 hours (Figure 3d; Table 1b; Figure 5). Visually, bubbles increased slightly in size and began to coalesce, suggesting a transition from predominantly nucleation-driven growth to a regime dominated by diffusive growth and coalescence.
Figure 4: Vertical profiles of gas saturation $S_{g}$ as a function of depth $z$ along the core axis at 8, 23 and 30 h. Depth increases downward from the top of the imaged sample and is expressed in micrometres; line colours correspond to the times indicated in the legend.
Vertical profiles of gas saturation $S_{g}(z)$ along the core axis (Figure 4) further illustrate this evolution. At 8 h the gas phase is confined to a thin zone near the top of the imaged interval, with $S_{g}$ close to zero over most of the depth range. By 23 h, gas saturations of $\sim 0.3$–0.8 extend through much of the interior of the sample, while only the basal few hundred micrometres remain largely brine saturated. The profiles retain a clear top bias, as expected for a low-density gas such as H2, but significant gas fractions are also present at depth, implying that capillary forces and local pore geometry are sufficient to trap and hold ganglia against buoyancy once nucleation has occurred. The profile at 30 h has a similar overall shape but slightly higher $S_{g}$ in parts of the column, consistent with modest late-time growth and coalescence of existing ganglia rather than widespread nucleation of new bubbles.
Overall, the time-lapse imaging and derived metrics support a three-stage evolution (Figures 3 and 5; Table 1b): (i) an initial fully saturated stage with no detectable gas (0–8 h), (ii) a rapid gas generation stage with strongly increasing $S_{g}$ (8–23 h), and (iii) a slower growth stage in which $S_{g}$ approaches an apparent plateau (23–30 h). Quantitative connectivity metrics (cluster statistics and percolation, including $f_{\max}$; Figure 5; Table 1b) further characterise changes in gas ganglia structure and will be discussed in Section 4.
Table 1: Summary of image-derived metrics used in the kinetic and connectivity analysis. $\phi$ is porosity, AA the solid–pore interfacial area within the imaged volume, $S_{g}$ the bulk gas saturation, $\Delta S_{g}/\Delta t$ the approximate rate of change between scans, and $f_{\max}$ the volume fraction of the largest gas cluster relative to total gas volume.
(a) Static micro-CT geometry at $t=0$.
$t$ $\phi$ $V_{s} AA$
(h) (-) ($\mathrm{m^{2}\,m^{-3}}$) ($\mathrm{m^{2}}$)
[…]
(b) Time-lapse gas saturation and connectivity metrics.
Scan $t$ $S_{g}$ $\Delta t$ $\Delta S_{g}/\Delta t$ $f_{\max}$ Perc.
(h) (-) (h) ($\mathrm{h^{-1}}$) (-) (Yes/No)
1 0 0.00 0 No
[…]
Largest cluster touches both the top and bottom boundaries of the imaged volume but does not form a fully continuous spanning path at the voxel scale.
Figure 5: Evolution of bulk gas saturation $S_{g}$ (left axis, solid blue line) and the fraction of gas in the largest cluster $f_{\max}$ (right axis, dashed orange line) as a function of time during the experiment. Shaded bands indicate the three stages inferred from the image-derived metrics: (i) no resolvable gas (0–8 h), (ii) rapid gas generation and connectivity growth (8–23 h), and (iii) slower, diffusion-limited growth (23–30 h).
Given the pre-removal of CO2 via the extensive 1000 PV brine flush, the absence of any intentionally introduced gas phase, and the fact that water remains well below its boiling point at 100C and 4 bar, the nucleating gas is most plausibly dominated by molecular H2 generated via hydrothermal alteration of the ultramafic grains. The high-contrast KI brine allows the gas phase to be segmented unambiguously against the rock matrix in the micro-CT images. Independent presence indicators further support this interpretation: during controlled depressurisation at the end of the experiment the electrochemical detector registered H2 concentrations up to $\sim$80 ppm in the pump gas effluent, and the H2-indicator strip in the confining sleeve showed a positive response. However, these measurements provide presence/threshold information rather than a full gas composition, so minor contributions from other volatiles (e.g., residual dissolved gases or water vapour) cannot be completely ruled out.
3.2 Mineralogical Alteration and Geochemistry
BEX composite images of the grain mounts show that both the pre- and post-reaction samples are compositionally heterogeneous at the grain scale, but the distribution and intensity of the colour-coded elemental response change after reaction (Fig. 6a,e). In the reacted sample, patches of higher intensity occur on a subset of grains in both the multi-element composite and the Fe-only maps (Fig. 6b,f), indicating localised changes in surface composition rather than a uniform alteration of all grains.
Post-experiment SEM imaging revealed distinct, localised textural changes compared to the unreacted control (Fig. 6c,d,g,h). In the pre-reaction sample, grain surfaces at the overview scale are relatively smooth and, at higher magnification, consist mainly of compact, equant to sub-rounded microcrystalline material with no obvious oriented growths (Fig. 6c,d). In contrast, the reacted grains show more irregular, roughened surfaces and, in places, clusters of elongate to platy crystals protruding from the grain surface into the pore space (Fig. 6g,h). These features occur in discrete patches rather than covering entire grains, and we interpret them as localised secondary alteration products.
Figure 6: Representative BEX and SEM images of the ultramafic grains before and after reaction. (a,e) Multi-element BEX composite images of pre- and post-reaction mounts showing compositionally heterogeneous but broadly Mg-rich grains, with locally enhanced signals after reaction. (b,f) Fe-only BEX maps highlighting Fe-enriched patches on a subset of grains. (c,g) SEM overview images of pre- and post-reaction grain surfaces. (d,h) Higher-magnification SEM zooms illustrating the transition from relatively smooth, compact pre-reaction surfaces (d) to roughened surfaces with discrete clusters of elongate to platy surface crystals protruding into the pore space in the reacted sample (h). These features are interpreted as localised secondary alteration products and compositional reorganisation at grain surfaces.
BEX spot analyses highlight corresponding local variations in elemental signals between the pre- and post-reaction grains (Table 2). Across the five pre-reaction spots, Mg- and Fe-rich compositions dominate and Al and Ca remain relatively low, whereas the post-reaction spots show a broader spread with locally elevated Al and Ca and Mg (and, to a lesser extent, Fe) extending to lower values. One plausible explanation is partial dissolution of Mg-rich primary phases, leaving behind or exposing domains that are comparatively enriched in Al and associated Ca-bearing material, possibly including thin secondary coatings where the elongate crystals are observed. We cannot completely exclude a contribution from salt precipitation upon drying, given the presence of KI in the brine, but in the areas analysed the BEX signals are dominated by Mg, Al, Ca and Fe and do not show elevated K or I consistent with pure KI crystals. Similar Al- and Ca-enriched alteration products have been reported in low-temperature alteration of ultramafic rocks and are commonly attributed to Al-bearing phyllosilicates or related secondary phases. However, the BEX data are semi-quantitative and each spot samples only a very small area; the observed ranges may therefore reflect a combination of reaction-driven changes and pre-existing micro-scale heterogeneity. We consequently treat the elevated Al and Ca signals as qualitative indicators of local mineralogical reorganisation at grain surfaces rather than as evidence for bulk-scale Al mobilisation or the formation of specific, well-identified Al-bearing mineral phases.
Table 2: Elemental composition from bulk XRF and BEX spot analyses before (pre) and after (post) reaction. BEX values give the minimum–maximum range of wt. % across five spot analyses on each mount. Dashes indicate elements not detected in any of the BEX spots.
Bulk XRF BEX pre BEX post
Element wt. % wt. % (range) wt. % (range)
O 48.10–51.92 49.48–51.58
C 5.09–6.64
Na 0.045
Mg 20.127 18.88–20.97 14.44–21.10
Al 2.213 1.44–2.50 3.09–6.58
Si 17.524 16.38–18.97 15.19–17.49
P 0.018
S 0.008
K 0.003 0.23–0.23
Ca 1.780 0.36–0.36 0.26–6.11
Ti 0.081 0.06–0.06
Mn 0.108 0.13–1.22 0.34–0.34
Fe 6.542 5.32–6.80 5.19–7.29
Ni 0.156 0.17–0.42 0.20–0.20
Cr 0.691 0.11–0.30 0.40–0.40
Sr 0.001
Zn 0.018
Ba 0.009
Zr 0.006
Yb 0.90–0.90
Taken together, the SEM-BEX observations are consistent with ongoing fluid–rock reaction in the vicinity of the pore space during the experiment, leading to modification of grain surfaces. A more complete characterisation of the alteration products and their compositions would require targeted phase identification (e.g., diffraction or spectroscopy) and more extensive microanalytical sampling, which are beyond the scope of this proof-of-concept study.
4 Discussion
The time series of pore-scale images and derived metrics shows three distinct regimes (Figures 3 and 5; Table 1): (i) an initial period with no resolvable gas phase (0–8 h), (ii) a rapid transition to a connected network of bubbles between 8 and 23 h, and (iii) a subsequent phase of more modest bubble growth between 23 and 30 h. Although our experiment was not designed for precise kinetic inversion, simple order-of-magnitude estimates, constrained by the image-derived porosity, surface area and gas saturations, help to show that these regimes are consistent with delayed supersaturation, nucleation and subsequent diffusion-limited growth of an H2 gas phase.
The delay before the first visible bubbles can be understood as the time needed to build up a sufficient concentration of dissolved H2 in the pore water. From the micro-CT field of view and core-holder geometry, the total sample volume is of order $V_{s}\sim 10^{-7}\,$\mathrm{m}^{3}$\$, and the segmented images give a porosity of $\phi=0.375$, so that the pore volume is $V_{p}\approx\phi V_{s}\approx 4\times 10^{-8}\,$\mathrm{m}^{3}$\$ ($\sim 0.04\,$\mathrm{m}\mathrm{L}$\$). Published measurements and correlations for H2 solubility in pure water and NaCl brines at temperatures near $100^{\circ}$C and H2 partial pressures of order 1–10 bar give equilibrium dissolved concentrations of $c_{\mathrm{sat}}\approx(2$$6)\times 10^{-4}\,$\mathrm{m}\mathrm{o}\mathrm{l}\,\mathrm{L}^{-1}$\$ (e.g., Wiesenburg1979_H2Solubility). At these conditions the pore fluid remains entirely in the liquid phase (water vapour pressure at 100C is $\sim$1 bar), so the free gas we observe cannot be attributed to boiling of the brine itself. For our pore volume this corresponds to a total amount of dissolved H2 at near-saturation of
$n_{\mathrm{sat}}\sim c_{\mathrm{sat}}V_{p}\approx(1\text{--}3)\times 10^{-8}\,$\mathrm{m}\mathrm{o}\mathrm{l}$\$, (1)
where the range reflects the spread in reported solubilities over the relevant temperature and pressure range rather than uncertainties in our sample geometry.
Batch experiments on olivine and partially serpentinized peridotite at similar temperatures report surface-normalised H2 generation rates spanning roughly $10^{-10}$$10^{-8}\,$\mathrm{m}\mathrm{o}\mathrm{l}\,\mathrm{m}^{-2}\,\mathrm{s}^{-1}$\$ depending on mineralogy and fluid composition (e.g., Mayhew2013; Neubeck2014; McCollom2016). Complementary flow-through experiments on olivine under hydrothermal conditions indicate comparable order-of-magnitude rates while highlighting the influence of fluid flow and evolving permeability on serpentinization and H2 release (Ross2025_OlivineFlow). The image-derived grain surface area accessible to the pore fluid within the imaged volume is $A\approx 4.3\times 10^{-4}\,$\mathrm{m}^{2}$\$ (Table 1a), so the implied total H2 production rate is
$R\sim rA\sim(4\times 10^{-14}\text{--}4\times 10^{-12})\,$\mathrm{m}\mathrm{o}\mathrm{l}\,\mathrm{s}^{-1}$\$. (2)
Combining this with the estimate of $n_{\mathrm{sat}}$ above gives a characteristic time to accumulate near-saturation concentrations of dissolved H2 of
$t_{\mathrm{sat}}\sim\frac{n_{\mathrm{sat}}}{R}\sim 2\times 10^{3}\text{--}7\times 10^{5}\,$\mathrm{s}$\$, (3)
i.e. roughly 0.50.5200200 hours when the full range of reported rates is used. If we instead adopt a mid-range literature value of $r\approx 10^{-9}\,$\mathrm{m}\mathrm{o}\mathrm{l}\,\mathrm{m}^{-2}\,\mathrm{s}^{-1}$\$, then $R\approx 4\times 10^{-13}\,$\mathrm{m}\mathrm{o}\mathrm{l}\,\mathrm{s}^{-1}$\$ and $t_{\mathrm{sat}}\approx(2$$7)\times 10^{4}\,$\mathrm{s}$\$, i.e. approximately 6–20 hours. This narrower range brackets, and is within a factor of a few of, the observed onset time of resolvable bubbles ($\sim$8 h). The close agreement between the simple solubility-based timescale and the measured induction period strongly supports an interpretation in which an initial phase of reaction-driven dissolved H2 accumulation is followed by gas-phase nucleation, rather than one in which bubbles are produced primarily by instantaneous thermal expansion of pre-existing air or exsolution of background dissolved gases during heating.
[…]
The rapid increase in gas occupancy between 8 and 23 h, from $S_{g}\approx 0.01$ to $S_{g}\approx 0.39$ (Table 1b), is consistent with a nucleation-and-growth phase in which many pores cross the supersaturation threshold over a relatively short time interval. In this regime, newly produced H2 partitions between further dissolution and growth of existing bubbles.
Assuming a diffusion coefficient for dissolved H2 in water of $D\sim 10^{-9}\,$\mathrm{m}^{2}\,\mathrm{s}^{-1}$\$, the diffusive length scale over a time $t$ is
$\ell_{D}\sim\sqrt{Dt}\approx 5\text{--}9\,$\mathrm{m}\mathrm{m}$\$ (4)
for $t$ between 8 and 23 h. This is comparable to the characteristic size of the imaged volume, implying that dissolved H2 can be redistributed throughout the sample and feed bubble growth over the timespan between scans.
By 23 h, the images show a large, internally connected network of gas ganglia occupying much of the pore space, suggesting that capillary entry pressures have been overcome in a significant fraction of the pore throats and that the system has approached a quasi-steady balance between continued H2 generation, dissolution and gas-phase growth.
The more modest changes observed between 23 and 30 h (Fig. 3c,d), during which $S_{g}$ increases only slightly from $\approx 0.39$ to $\approx 0.43$ and the apparent rate $\Delta S_{g}/\Delta t$ drops by almost an order of magnitude (Table 1b), are naturally interpreted as a late-time regime in which (i) the driving force for further H2 exsolution is reduced as dissolved concentrations approach a new equilibrium with the gas phase, (ii) reactive surface area is locally reduced as alteration rinds and secondary phases develop on grain surfaces, and (iii) additional H2 produced at depth can diffuse and be taken up by existing bubbles without requiring widespread nucleation of new gas domains. In this view, the observed time sequence reflects a transition from a reaction-controlled build-up of dissolved H2 (pre-nucleation), through a short nucleation-and-growth episode once supersaturation is reached, to a diffusion- and surface-reaction–limited regime at later times.
Although these back-of-the-envelope calculations are necessarily approximate, they demonstrate that the qualitative evolution seen in the micro-CT time series—now quantified in terms of measured porosity, surface area and gas saturation—is consistent with known H2 production rates during low-temperature alteration of ultramafic rocks and with simple transport and solubility constraints. More broadly, the results highlight that even in a small, static system, there can be a substantial lag between the onset of water–rock reaction and the formation of a resolvable free gas phase, followed by a relatively short period during which pore-scale connectivity of the gas is established. Such behaviour has important implications for natural H2 systems: it suggests that zones of intense reaction may remain “invisible” in terms of free gas for extended periods, and that the conditions required to cross nucleation and connectivity thresholds may be highly localised in space and time.
We did not perform a separate brine-only micro-CT control experiment in an inert granular medium, but the early part of our time series effectively provides an internal control: the same apparatus, KI brine and loading conditions show no resolvable gas for the first 8 h, with a gas phase only appearing after prolonged reaction. This behaviour is difficult to reconcile with a purely instrumental or brine-related artefact and instead points to the role of water–rock reaction in driving gas exsolution.
4.1 Implications for natural H2 systems
Although this experiment is deliberately small-scale and exploratory, the observed sequence of (i) delayed bubble appearance, (ii) rapid establishment of a connected gas network, and (iii) later-time, slower growth has several implications for natural H2 systems and engineered analogues.
First, the long pre-nucleation period implies that substantial water–rock reaction and H2 production can occur before a free gas phase becomes detectable at the pore scale. In natural settings, zones of active H2 generation may therefore be dominated by dissolved H2 for extended periods, with little or no expression as free gas until local supersaturation and favourable nucleation conditions are met. Experimental work on fractured ultramafic rocks likewise shows that serpentinization can rapidly seal existing flow paths by reducing fracture permeability by orders of magnitude, implying that sustained fluid and gas transport requires continual generation of new permeability (Farough2016). This behaviour is directly relevant to proposals for exploiting rock-based hydrogen resources, including “orange hydrogen” schemes that couple water–rock–H2 generation with carbon mineralisation and storage (Osselin2022).
Second, once nucleation thresholds are crossed, the rapid transition to a connected gas ganglia suggests that gas connectivity may be established over relatively short timescales compared to the overall reaction history. This is consistent with the view that small changes in capillary entry pressures or wettability can trigger large changes in effective gas mobility (e.g., Thaysen2023; Jangda2023; Rezaei2022). In heterogeneous natural reservoirs, such threshold behaviour is likely to be highly localised, so migration pathways for natural H2 may be controlled by a limited number of “activated” reaction zones.
Third, the modest changes between 23 and 30 h highlight that, once a gas phase is established, further H2 production can be partitioned between dissolution into the surrounding fluid and incremental growth of existing bubbles rather than continued nucleation. This behaviour is directly relevant to both natural H2 accumulations and underground H2 storage, where the balance between dissolution, trapping and mobile gas strongly influences storage security and recoverable volumes (Heinemann2021; Miocic2023). Our results emphasise that predicting this balance requires not only bulk reaction kinetics, but also an understanding of pore-scale nucleation thresholds and the evolving microstructure of reactive rocks.
Overall, the experiment demonstrates that even in a simple, static system, bubble nucleation and gas connectivity are controlled by the interplay of reaction, transport, and capillary processes. Extending this type of pore-scale, time-resolved imaging to a wider range of mineralogies, p–T conditions and flow regimes will be essential for translating geochemical models of H2 generation into quantitative predictions of where and when a free H2 phase can form, migrate and accumulate in the subsurface.
5 Conclusions
This proof-of-concept experiment demonstrates that time-resolved 4D micro-CT can be used to monitor the emergence and early evolution of a gas phase generated by water–rock reaction in an ultramafic granular pack at 100C. Because the results are derived from a single experiment on a crushed granular pack under one set of conditions, they should be viewed as qualitative constraints on pore-scale behaviour rather than quantitative predictions for specific field settings. After an initial period of $\sim$8 hours with no resolvable gas, bubbles nucleate and rapidly form a connected ganglia network, followed by a slower growth phase. Simple order-of-magnitude estimates of H2 solubility, reaction rates and diffusion lengths show that this three-stage evolution is consistent with delayed supersaturation, threshold-controlled nucleation and subsequent diffusion-limited growth within the imaged volume, and the predicted saturation timescale (6–20 hours) closely matches the observed induction period. This agreement provides strong support for the interpretation that the gas is predominantly reaction-derived H2 rather than air or steam.
Independent gas measurements and H2-indicator strips are consistent with the imaged gas phase being predominantly H2, although the experiment was not designed to uniquely determine gas composition or to fully resolve reaction pathways. SEM-BEX observations provide qualitative evidence for localised surface alteration and the development or exposure of secondary phases on grain surfaces, but they do not by themselves demonstrate bulk Al mobilisation or allow unambiguous phase identification. We therefore interpret both the gas and mineralogical observations as qualitative indicators of ongoing fluid–rock reaction and associated gas generation in the sample.
Taken together, the results highlight that a significant amount of low-temperature water–rock reaction can occur before a free gas phase becomes detectable at the pore scale, and that once nucleation thresholds are crossed, gas connectivity can be established over relatively short timescales compared to the overall reaction history. This behaviour has direct implications for natural H2 systems and underground H2 storage, where the balance between dissolved, trapped and mobile gas phases controls migration and recoverability. Future work should extend this approach to a broader range of mineralogies, pressures, temperatures and flow regimes, and combine pore-scale imaging with more detailed geochemical and mineralogical characterisation to better quantify the coupling between serpentinization kinetics, gas generation and multiphase transport in reactive ultramafic rocks. The primary contribution of this study is therefore the direct, spatially and temporally resolved observation of gas nucleation and connectivity during low-temperature alteration, rather than detailed identification of the neoformed mineral phases. Future work combining similar imaging with complementary mineralogical and geochemical analyses will be essential to fully resolve the reaction pathways involved.
Open Research Section
The micro-CT image volumes, processed segmentations, and representative SEM-BEX images used in this study will be made available in an open-access repository (e.g., Zenodo) upon acceptance of the manuscript. A persistent DOI and metadata will be provided in the final published version.
Conflict of Interest declaration
The authors declare there are no conflicts of interest for this manuscript.
Acknowledgements.
This research was funded by The Carnegie Trust RIG013387. Sample collection was funded by the Gilchrist Educational Trust (grant GILC 06.22, administered by the Royal Geographical Society).
Author Contributions
H.P.M and A.G. conceptualised this study. H.P.M. raised funding. A.G. chose the sample. M.W. prepared the sample. H.P.M. and Z.Z.J. did the imaging experiment. J.B. did the pre and post experimental SEM-BEX imaging. Z.Z.J. and H.P.M. analysed the micro-CT image data. M.W. analysed the SEM-BEX data. H.P.M. wrote the first draft. All authors contributed to writing this paper.